on October 1 Amy inc purchased equipment for 40000 and estimates monthly depreciation of $500. What will be the book value of the building on December 31

a)40,000
1,500
1000
d) 39,000
e) 38500
Answer #1

e) 38,500

Number of months covered as on dec 31= 3 months

Total depreciation for 3 months= 3*500= $1,500

Bookvalue as on Dec 31= 40,000 - 1,500

= 38,500